The unit of ratio of magnetic field and electric field is?

Grade:12

1 Answers

Gulshan raj
7 years ago
Lorentz force =qE +qvB Considering E is perpendicular to B. Then electric force and magnetic force cancel each other remaining with F =0.Hence,, qE =qvB =>V = E/BV= velocity E=electric field and B =magnetic field. Unit of ratio of E/B is m/secAnd ratio of B/E is sec /m.